If 0.65 moles of hydrogen occupies 9.0 L, how many moles must be added to inflate the balloon to 15.6 L?(round to two sig figs)


1
Expert's answer
2021-08-19T04:41:04-0400

n1 = 0.65 mol

V1 = 9.0 L

n2 = unknown

V2 = 15.6 L


Solution:

PV = nRT

V/n = RT/P

RT/P = const

Therefore,

V1/n1 = V2/n2

n2 = n1V2 / V1

n2 = (0.65 mol × 15.6 L) / (9.0 L) = 1.1267 mol = 1.1 mol (two sig figs)

n2 = 1.1 mol


Answer: 1.1 mol of hydrogen gas must be added to inflate the balloon to 15.6 L.